OpenStudy (anonymous):

write the equation of the circle in standard form. x^2+y^2-10x-4y+13=0. First, i brought 13 over to the left side of the equation, and then grouped terms. (x^2-10x+___)+(y^2-4y+___)=-13 Then, i completed the square by dividing the numbers before the x's in half, and then squaring them to get... (x^2-10x+25)+(y^2-4y+4)=-13+25+4 Simplify... (x-5)^2+(y-2)^2=16 at this point, your equation is in standard form. r^2=16,so r=4 the center is at (h,k) which is (5,2) and your intercepts are going to be when you set x=0 and y=0 and solve for each. For y, -5+(y-2)=4 y-2=9 y=11 (0,11)x=(7,0). Something is wrong! help.

jimthompson5910 (jim_thompson5910):

your center and radius are correct, so nice work on getting that

jimthompson5910 (jim_thompson5910):

x-intercepts: (x-5)^2+(y-2)^2=16 (x-5)^2+(0-2)^2=16 (x-5)^2+(-2)^2=16 (x-5)^2+4=16 (x-5)^2=16-4 (x-5)^2=12 x-5 = sqrt(12) or x-5 = -sqrt(12) x-5 = 2*sqrt(3) or x-5 = -2*sqrt(3) x = 5 + 2*sqrt(3) or x = 5 - 2*sqrt(3) Those are your exact x-intercepts. You would use a calculator to approximate them

jimthompson5910 (jim_thompson5910):

y-intercepts (x-5)^2+(y-2)^2=16 (0-5)^2+(y-2)^2=16 25 + (y-2)^2 = 16 (y-2)^2 = 16 - 25 (y-2)^2 = -9 y - 2 = sqrt(-9) or y - 2 = -sqrt(-9) since the square root of a negative number doesn't produce a real number, this means that there are no real solutions to (0-5)^2+(y-2)^2=16 so there are no y-intercepts

jimthompson5910 (jim_thompson5910):

here is the graph of the circle

OpenStudy (anonymous):

thank you so much! you're awesome.

jimthompson5910 (jim_thompson5910):

you're welcome